A train of 18 carriages each of 50m length with an engine of 50m length is running at 50kmph. find the time this train crosses a bridge of 2 km length
Answer by lwsshak3(11628)

A train of 18 carriages each of 50m length with an engine of 50m length is running at 50kmph. find the time this train crosses a bridge of 2 km length
***
travel time=distance/speed
distance=18*50+50+2000=2950=2.95 km
travel time=2.95/50=0.059
time train crosses bridge=0.059 hr or 3.45 min